OpenStudy (anonymous):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura returned the movie and paid $5.00, how many nights did she keep the movie?

OpenStudy (goldphenoix):

You can use an equation to solve this. The flat fee is constant, which mean that it cannot change. .50 dollar is not constant. So the equation should look like: \[2.00 + 0.50x=5.00\] To find X, you want to get rid of 2. To get rid of 2, you must subtract 2 from both side. 2 + .50x = 5 -2 -2 ----------------- .50x=3. The final step is to divide both side by .50. \[\frac{ 3 }{ .5 } = 6\] So x =6. X mean the amount of nights.